A scientist inoculates several mice, one at a time, with a disease germ until he finds 3 mice that have attacked by the disease. If the probability of getting attack of the disease is 1/5, what is the probability that 10 mice are required


1
Expert's answer
2022-05-17T15:22:30-0400

Apply the Negative Binomial distribution.

Let "n=10" be the sample size of the sample until we get "r=3" successes. The probability of contracting the disease is "p=1\/5.". The form of the Negative Binomial distribution we apply is given as,


"p(X=n)=\binom{n-1}{r-1}p^r(1-p)^{n-r}"

.Therefore,


"p(X=10)=\binom{10-1}{3-1}(\dfrac{1}{5})^3(1-\dfrac{1}{5})^{10-3}"

"=36(0.008)(0.2097152)=0.0603979776"